OpenStudy (anonymous):

A square pyramid has a slant height of 13 meters and a height of 12 meters. What is the surface area if the square base has sides of length 10 meters? 360 m2 320 m2 260 m2 120 m2

OpenStudy (experimentx):

4*1/2*slant height*base + base * base

OpenStudy (anonymous):

so confused

Directrix (directrix):

The surface area of a regular pyramid is equal to the area of its base plus 1/2 times the perimeter of its base times the slant height. SA = 10*10 + (1/2)*(40)(13) SA = 100 + 260 SA = 360 square meters

Directrix (directrix):

Questions?

OpenStudy (anonymous):

ok that makes more sence